OpenStudy (anonymous):

Penny uses 3/4 yard of fabric for each pillow she makes. how many pillow can she make using 6 yards of fabric? Do I turn 6 into a fraction, use the reciprocal and cross multiply?

OpenStudy (anonymous):

first off, there is no such thing as "cross multiply"

OpenStudy (anonymous):

In my school they use cross multiplying as simplify im only in 6th grade

OpenStudy (anonymous):

divie \(6\) by \(\frac{2}{4}\) which is the same as multiplying \(6\) times \(\frac{4}{3}\)

OpenStudy (anonymous):

tell your teacher satellite73 said there is no such mathematical operation as "cross multiply" and to remove that from his or her vocabulary

OpenStudy (anonymous):

answer is 7 and 1/2 pillow cases

OpenStudy (anonymous):

I sure will!

OpenStudy (anonymous):

in any case your job is to compute \[\large 6\div \frac{3}{4}=6\times \frac{4}{3}\]

OpenStudy (anonymous):

its called cross simplifying

OpenStudy (anonymous):

cancel first, multiply last the answer is not seven and a half

OpenStudy (anonymous):

there is not such thing as "cross simplifying" either or "simplify" for that matter

OpenStudy (anonymous):

Honestly rounding that was mental math and it is close to the answer and it is cross products

OpenStudy (anonymous):

I got 8 as an answer, is it correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

three goes in to six two times, two times four is eight

OpenStudy (anonymous):

Ever heard of cross products...

OpenStudy (anonymous):

I simplified both fractions

OpenStudy (anonymous):

yes of course for vectors

OpenStudy (anonymous):

Ps i might sound dum because im really really really tired

OpenStudy (anonymous):

6/1 / 3/4 = 6/1 * 4/3 = 24/3 = 8 so 8 pillowcases

OpenStudy (anonymous):

And cross multiplying is a term used in Core Connections/ CPM 6th and 7th grade math.

OpenStudy (anonymous):

And i am not in a mental state of making right choices because I just took a shower which "Woke" me up and I can't fall asleep

OpenStudy (anonymous):

ima go help people with science :)#Science is the best!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

It is basically what i did above just taken out of fraction form: